### v2

# Relative Pose Estimation for Nonholonomic Robot Formation with UWB-IO Measurements

Abstract
This article studies the problem of distributed formation control for multiple robots by using onboard ultra wide band (UWB) distance and inertial odometer (IO) measurements. Although this problem has been widely studied, a fundamental limitation of most works is that they require each robot’s pose and sensor measurements are expressed in a common reference frame. However, it is inapplicable for nonholonomic robot formations due to the practical difficulty of aligning IO measurements of individual robot in a common frame. To address this problem, firstly, a concurrent-learning based estimator is firstly proposed to achieve relative localization between neighboring robots in a local frame. Different from most relative localization methods in a global frame, both relative position and orientation in a local frame are estimated with only UWB ranging and IO measurements. Secondly, to deal with information loss caused by directed communication topology, a cooperative localization algorithm is introduced to estimate the relative pose to the leader robot. Thirdly, based on the theoretical results on relative pose estimation, a distributed formation tracking controller is proposed for nonholonomic robots. Both 3D and 2D real-world experiments conducted on aerial robots and grounded robots are provided to demonstrate the effectiveness of the proposed method.
Index Terms:
Relative localization, nonholonomic robots, local reference frame, formation control.
I Introduction
Distributed formation control for robot swarms has received significant attention in recent years due to its wide range of potential applications [1]. The objective of formation control is to guide robot swarms from an initial configuration to a specified formation through distributed interactions. To achieve this objective, various methods have been explored, including those based on consensus [2], artificial light field [3] and mean-shift exploration [4]. However, most of these methods rely on external localization systems, such as GPS, which are not well-suited for practical applications. Moreover, external localization systems may be unavailable in certain scenarios, such as indoor or underground environments. In summary, compared to relying on external localization systems, robots that depend solely on onboard measurements are more autonomous.
Extensive researches adopted sensor measurements to estimate relative poses between neighboring robots. Based on the type of sensor utilized, formation control methods relying on local measurements can be roughly classified into two categories: relative position, (bearing or angle-based) methods, and distance-based methods. For the first class methods, depth camera or LADAR can be used to achieve ralative measurements between neighboring robots. Distributed formation control task can be accomplished when the inter-robot measurement topology satisfies specific conditions such as bearing rigidity [5] or non-collinear neighors condition [6, 7]. To relax the topology condition, recent works [8, 9] enhance the observability of relative pose of neighbor robots by more effectively utilizing visual measurement information. However, all these methods require expensive and complex sensors and object detection algorithms, significantly increasing both complexity and hardware costs. Additionally, these methods face challenges like limited field of view and high environmental requirements, including lighting conditions [10].
Figure 1: Geometric relationship between the displacement and UWB distance measurements of the two robots. A. All the robot have a global orientation [11, 12, 13, 14, 15, 16, 17]. B. Each robot is in its local frame.
Distance measurement sensors like Ultra Wide Band (UWB) inherently address the challenges [18, 19]. One of the primary difficulties of formation control that relies on distance measurements is accurately estimating. To estimate relative position, traditional methods typically derive the leader’s relative position using distance measurements from multiple neighbors [20]. However, this method inevitably imposes conditional constraints on the measurement topology, such as the requirement of rigidity. Similarly, a common strategy is to place multiple UWB tags on each robot [21, 22, 23] or install UWB anchors in the environment [24], which may sacrifice the scalability of the robot swarm. Another strategy involves introducing the inertial odometer (IO) measurement to achieve relative localization between robots [11]. Using a relative position estimator based on UWB distance and local IO measurements, distributed formation control of robot swarms has been achieved [12, 13, 14, 25]. To handle the sensor characteristics, the IO scale factor is estimated in [15], and an abnormal data removal mechanism is proposed in [16].
All the aforementioned methods address the inter-robot relative localization problem within a global reference frame, as illustrated in Fig. 1 A. However, for nonholonomic robots, relative pose estimation between robots must be conducted in a distributed manner within the local coordinate frame, as shown in Fig. 1 B. This poses significant challenges when relying solely on relative distance measurements. Fortunately, by increasing the IO displacement of each robot within its odometer frame, the relative yaw angle between the two robots can be estimated [26, 27, 10]. The core concept involves formulating an optimization problem that relates real-time distance measurements and the relative pose between robots. In these methods, only UWB distance and local IO measurements are adopted to estimate relative pose estimation between two robots. Nevertheless, all these methods [12, 14, 27, 10] require that the relative motion between robots satisfy the persistently excited (PE) condition to ensure the convergence of the relative localization estimators. In other words, these methods require that robots maintain a non-zero relative velocity at all times. This is typically achieved through the introduction of continuous excitation noise [14] and periodic oscillation signals [11, 10] of varying amplitudes into the control inputs of the robots. The disadvantage lies in the contradiction between the motion needed for PE localization enhancement and relatively stationary formation control tasks, which can result in uneven motion trajectories and diminish control performance. To address this issue, a rotating UWB tag has been used to achieve the PE condition [17]. However, this solution significantly increases the complexity of the sensor equipment on each robot and raises the failure rate. So far, there are few studies focused on the implementation of distributed formation control for multiple nonholonomic robots using only UWB and IO measurements.
The primary main challenge lies in achieving distributed cooperative pose estimation based on relative distance measurements without necessitating the PE condition for robot relative motion. This article aims to address the above challenge, and the contributions can be summarized as follows. First, the proposed distributed formation control law relies solely on inter-robot distance measurements. Unlike existing formation control methods for nonholonomic robots that require knowledge of inter-robot relative positions [1, 7, 28] or relative orientations [7, 29], our method utilizes only UWB distance and IO measurements. This reduces the sensor requirements for each robot and broadens applicability across various scenarios. Second, to facilitate relative pose estimation between robots in a local frame, a concurrent-learning based relative pose estimator is proposed. Different from recent methods[12, 14, 1] that estimate relative position or orientation in a global frame, our approach focuses on estimating relative pose in a local context. In contrast to PE relative motion based methods [14, 12, 27], our method effectively incorporates both current and historical information to relax the PE condition to a sufficiently excited (SE) condition in the relative pose estimation problem. This can effectively improve the smoothness of robot trajectories and enhance control performance. Additionally, the proposed data-driven methods demonstrate superior noise robustness compared to those relying solely on current information [14]. We provide both theoretical analysis and real-world experiments involving mobile robots and drones are provided to validate the effectiveness of the proposed methods.
The rest of this article is organized as follows. In Section II, the distributed formation control task for nonholonomic robots is formulated. In Section III, the main results for inter-robot relative pose estimation, leader relative pose estimator design and the distributed formation controller are introduced, followed by the performance evaluation and experimental results in Section IV. Finally, Section V concludes this article.
Notations: A transformation matrix in frame $\{A\}$ is denoted as $\textbf{T}^{A}:=\begin{bmatrix}\textbf{R}^{A},&\textbf{p}^{A}\\ \textbf{0}^{T},&1\\ \end{bmatrix}\in SE(3)$, where $\textbf{p}^{A}\in\mathcal{R}^{3}$ and $\textbf{R}^{A}\in SO(3)$ are the position vector and rotation matrix in frame $\{A\}$, respectively. Denote $\textbf{T}^{A}_{B}$ and $\textbf{R}^{A}_{B}$ as the transformation and rotation matrices from frame $\{B\}$ to frame $\{A\}$, respectively. For vector x, $x[i]$ and $x[i:j]$ denote the $i$-th element and the elements from $i$ to $j$ of x. For ease of reading, we list the main symbols of this article in Appendix A-A.
II Preliminaries and problem formulation
II-A Robot Models
Consider a group of $N+1$ nonholonomic robots in $\mathbb{R}^{2}$ or $\mathbb{R}^{3}$. For $i=0,...,N$, the kinematics of robot $i$ is described as
$\dot{\textbf{p}}^{\mathcal{O}_{i}}_{i}(t)&=\begin{bmatrix}\cos\phi_{\Sigma_{i}}^{\mathcal{O}_{i}}(t),&0\\ \sin\phi_{\Sigma_{i}}^{\mathcal{O}_{i}}(t),&0\\ 0,&1\end{bmatrix}\begin{bmatrix}v_{i,h}\\ v_{i,z}\end{bmatrix}\\ \dot{\phi}_{\Sigma_{i}}^{\mathcal{O}_{i}}(t)&=w_{i}.\end{aligned}\right.$ (1)
For each robot $i$, denote the odometry frame as $\mathcal{O}_{i}$. The $x$-axis of $\mathcal{O}_{i}$ is the initial orientation of robot $i$, as shown in Fig. 1 B. At time $t$, denote the IMU body frame as $\Sigma_{i}(t)$. Similarly, the $x$-axis of $\Sigma_{i}(t)$ is the defined as the real-time orientation of robot $i$. $\textbf{p}^{\mathcal{O}_{i}}_{i}(t)$ is the displacement in its odometry frame $\mathcal{O}_{i}$ from $t_{0}=0$ to $t$, which can be measurement with IO. $v_{i,h}$ and $v_{i,z}$ are the horizontal and vertical velocities of robot $i$. Similarly, $\phi_{\Sigma_{i}}^{\mathcal{O}_{i}}(t)$ is the angular displacements of robot $i$ in its odometry frame $\mathcal{O}_{i}$ at time $t$, which is also measurable by IO.
Note that the roll and pitch angles of the robot can be accurately measured with the IMU-based SLAM methods [10]. We focus on the 4-DoF robot described by Eq.(1), which can well describe both 2D and 3D nonholonomic robots.
II-B UWB Measurement and Communication Topology
Each robot $i$ is equipped with a UWB module to measure the real-time distance and communicate with each other. Define the position of robot $j$ in frame $\mathcal{O}_{i}$ at time $t$ as $\textbf{p}^{\mathcal{O}_{i}}_{j}(t)$. Then the relative position between robot $i$ and robot $j$ in frame $\mathcal{O}_{i}$ can be denoted as $\textbf{p}^{\mathcal{O}_{i}}_{ij}(t)=\textbf{p}^{\mathcal{O}_{i}}_{i}(t)-\textbf{p}^{\mathcal{O}_{i}}_{j}(t)$. With the UWB module, each robot $i$ can measure the real-time distance $d_{ij}(t)\triangleq\lVert\textbf{p}^{\mathcal{O}_{i}}_{ij}(t)\rVert$ from itself to its neighbor $j$ and communicate with each other. The interaction network among the robots is described by a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{0,1,\ldots,N\}$ and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$. Robot 00 is the leader robot. If $(i,j)\in\mathcal{E}$, robot $i$ and $j$ are neighbors, neighbors can communicate and measure distance with each other. The neighbor set of robot $i$ is denoted as $\mathcal{N}_{i}=\{j\in\mathcal{V}:(i,j)\in\mathcal{E}\}$.
II-C Control Objective
As shown in Fig. 1, under the condition that there are no global coordinate system and a shared orientation, the initial position and orientation of each robot $i$ is random. For each robot $i$, only the distance $d_{ij}$ to its neighbor $j$, odometer displacements $\textbf{p}^{\mathcal{O}_{i}}_{i}(t)$ and $\phi_{\Sigma_{i}}^{\mathcal{O}_{i}}(t)$ are measurable. Define the real-time relative orientation between robot $j$ and robot $i$ as $\theta_{\Sigma_{i}}^{\Sigma_{0}}(t)$. The control objective can be formulated as
$&\lim_{t\to\infty}\theta_{\Sigma_{i}}^{\Sigma_{0}}(t)=0\\ &\lim_{t\to\infty}\textbf{p}^{\mathcal{O}_{0}}_{i0}(t)-\textbf{p}_{i}^{*}=0,\end{aligned}\right.$ (2)
where $\textbf{p}_{i}^{*}\triangleq[p_{ix}^{*},p_{iy}^{*},p_{iz}^{*}]^{T}$ denotes the desired relative position for robot $i$ to the leader. Eq. (2) indicates that each robot $i$ has the same orientation and maintains a desired relative position $\textbf{p}_{i}^{*}$ with the leader in the leader’s odometer frame $\mathcal{O}_{0}$. To achieve the control task, the following lemma is required,
Lemma 1.
[1] Consider a perturbed system
$\dot{\textbf{x}}=f(t,\textbf{x})+g(t,\textbf{x},\bm{\delta}).$ (3)
where $\textbf{x}\in\mathbb{R}^{n_{x}}$ is the state and $\delta\in\mathbb{R}^{n_{\delta}}$ is an exogenous signal. $f(t,\textbf{x})$ and $g(t,\textbf{x},\bm{\delta})$ are piecewise continues in $t$ and locally Lipschitz in x and $(\textbf{x},\bm{\delta})$, respectively. Normal system $=f(t,\textbf{x})$ x ) in (3) is uniformly exponentially stable and $g(t,\textbf{x},\bm{\delta})$ satisfies $\lVert g(t,\textbf{x},\bm{\delta})\rVert\leq a_{1}\lVert\bm{\delta}\rVert\lVert\textbf{x}\rVert+a_{2}\lVert\bm{\delta}\rVert$, with two positive constants $a_{1}$ and $a_{2}$. Then system (3) uniformly asymptotically stable when $\lVert\bm{\delta}\rVert$ converges to 0 asymptotically.
It is worth noting that absolute localization systems, such as GPS or external global anchor points, are not necessary for accomplishing the formation control task. The leader robot provides a common reference frame for the swarm using its odometer frame. Other robots achieve formation tasks by collaboratively estimating their relative pose to the leader. If the current leader is damaged, the swarm can simply select a new leader, and the other robots can continue to perform cooperative localization to the new leader using our method.
III Methods
To achieve the control task, we design a strategy for each robot $i$ including neighbor relative localization, leader cooperative localization and local formation tracking controller. Firstly, we note that the real-time relative pose estimation problem between robot $i$ and robot $j$ can be formulated as the initial relative pose $\textbf{T}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}$ estimation problem (see Section III-A). The second step involves designing concurrent-learning based estimators $\hat{\textbf{p}}^{\mathcal{O}_{i}}_{ij,0}$ and $\hat{\textbf{R}}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}$ to estimate the initial relative pose between neighboring robots (see Section III-B). However, not all the robots can estimate its relative pose to the leader robot directly. To handle this issue, the third step is to design the estimators $\hat{\textbf{q}}^{\mathcal{O}_{i}}_{i0,0}$ and $\hat{\textbf{Q}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}$ to estimate the initial relative pose to the leader robot (see Section III-C). Finally, based on the relative pose estimators and odometer information. distributed formation controller $v_{i}$ and $w_{i}$ are devised (see Section III-D).
III-A Problem Formulation of Relative Localization
First, we show that relative localization between robot $i$ and its neighbor $j\in\mathcal{N}_{i}$ is equal to estimate the following relative frame transformation in robot $i$’s odometry frame $\mathcal{O}_{i}$:
$\textbf{T}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}:=\begin{bmatrix}\textbf{R}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}},&\textbf{p}^{\mathcal{O}_{i}}_{ij}(t_{0})\\ \textbf{0}^{T},&1\end{bmatrix},$
where $\textbf{p}^{\mathcal{O}_{i}}_{ij}(t_{0})$ is the initial relative position in frame $\mathcal{O}_{i}$. Define $\theta_{\mathcal{O}_{j}}^{\mathcal{O}_{i}}(t_{0})$ as the initial relative orientation between frame $\mathcal{O}_{j}$ and frame $\mathcal{O}_{i}$ at $t_{0}$. The rotation matrix from frame $\mathcal{O}_{j}$ to frame $\mathcal{O}_{i}$ can be defined as
$\textbf{R}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}:\triangleq\begin{bmatrix}\textbf{R}^{\mathcal{O}_{i,h}}_{\mathcal{O}_{j,h}},&\textbf{0}\\ \textbf{0},&\textbf{1}\end{bmatrix}=\begin{bmatrix}\cos\theta_{\mathcal{O}_{j}}^{\mathcal{O}_{i}}(t_{0}),&-\sin\theta_{\mathcal{O}_{j}}^{\mathcal{O}_{i}}(t_{0}),&0\\ \sin\theta_{\mathcal{O}_{j}}^{\mathcal{O}_{i}}(t_{0}),&\cos\theta_{\mathcal{O}_{j}}^{\mathcal{O}_{i}}(t_{0}),&0\\ 0,&0,&1\end{bmatrix}.$
Define the relative orientation between frame $\Sigma_{j}$ in frame $\Sigma_{i}$ at time $t$ as $\theta_{\Sigma_{j}}^{\Sigma_{i}}(t)$. With the IO measurements, the following equations holds,
$\textbf{p}^{\mathcal{O}_{i}}_{ij}(t)$ $=\textbf{p}^{\mathcal{O}_{i}}_{ij}(t_{0})+\textbf{p}^{\mathcal{O}_{i}}_{i}(t)-\textbf{R}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}\textbf{p}^{\mathcal{O}_{j}}_{j}(t)$ (4a)
$\theta_{\Sigma_{j}}^{\Sigma_{i}}(t)$ $=\theta_{\mathcal{O}_{j}}^{\mathcal{O}_{i}}(t_{0})+\phi_{\Sigma_{i}}^{\mathcal{O}_{i}}(t)-\phi_{\Sigma_{j}}^{\mathcal{O}_{j}}(t),$ (4b)
One can see that with IO measurements, $\textbf{p}^{\mathcal{O}_{i}}_{i}(t)\triangleq[\textbf{p}^{\mathcal{O}_{i}}_{i,h}(t),p^{\mathcal{O}_{i}}_{i,z}(t)]^{t}\triangleq[p^{\mathcal{O}_{i}}_{i,x}(t),p^{\mathcal{O}_{i}}_{i,y}(t),p^{\mathcal{O}_{i}}_{i,z}(t)]^{T}$ and $\phi_{\Sigma_{i}}^{\mathcal{O}_{i}}(t)$ are available. Furthermore, the neighbor’s IO measurements $\textbf{p}^{\mathcal{O}_{j}}_{j}(t)$ and $\phi_{\Sigma_{j}}^{\mathcal{O}_{j}}(t)$ can also be obtainable through UWB module. As a result, the relative localization problem between neighboring robots is equal to estimate the relative frame transformation $\textbf{T}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}$ between odometry frames $\mathcal{O}_{i}$ and $\mathcal{O}_{j}$.
At sampling time $t_{k}\triangleq k\Delta t,k\in\mathbb{N}$, the distance between robot $i$ and robot $j$ can be measured with the UWB module, as shown in Fig. 1 B. According to the cosine theorem, the measurements obtained from two adjacent samples satisfy
$\textbf{u}^{\mathcal{O}_{i}}_{ij}(t_{k})^{T}\textbf{p}^{\mathcal{O}_{i}}_{ij}(t_{k})=$ $\frac{1}{2}\Big{(}d_{ij}(t_{k+1})^{2}+d_{ij}(t_{k})^{2}-$ (5)
$\textbf{u}^{\mathcal{O}_{i}}_{ij}(t_{k})^{T}\textbf{u}^{\mathcal{O}_{i}}_{ij}(t_{k})\Big{)},$
where $\textbf{u}^{\mathcal{O}_{i}}_{ij}(t_{k})$ is the relative displacement in frame $\mathcal{O}_{i}$ from time $t_{k}$ to $t_{k+1}$ and it is defined as
$\textbf{u}^{\mathcal{O}_{i}}_{ij}(t_{k})=\textbf{u}^{\mathcal{O}_{i}}_{i}(t_{k})-\textbf{R}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}\textbf{u}^{\mathcal{O}_{j}}_{j}(t_{k}),$ (6)
where $\textbf{u}^{\mathcal{O}_{i}}_{i}(t_{k})=\textbf{p}^{\mathcal{O}_{i}}_{i}(t_{k+1})-\textbf{p}^{\mathcal{O}_{i}}_{i}(t_{k})\triangleq[\textbf{u}^{\mathcal{O}_{i}}_{i,h}(t_{k}),u^{\mathcal{O}_{i}}_{i,z}(t_{k})]^{T}\triangleq[u^{\mathcal{O}_{i}}_{i,x}(t_{k}),u^{\mathcal{O}_{i}}_{i,y}(t_{k}),u^{\mathcal{O}_{i}}_{i,z}(t_{k})]^{T}$ is the displacement of robot $i$ in frame $\mathcal{O}_{i}$ from $t_{k}$ to $t_{k+1}$, similar to $\textbf{u}^{\mathcal{O}_{j}}_{j}(t_{k})$. Substituting (4), (6) into (5), yields
$\textbf{u}^{\mathcal{O}_{i}}_{i}(t_{k})^{T}\textbf{p}^{\mathcal{O}_{i}}_{ij}(t_{0})+\textbf{u}^{\mathcal{O}_{i}}_{i}(t_{k})^{T}\textbf{p}^{\mathcal{O}_{i}}_{i}(t_{k})-\textbf{u}^{\mathcal{O}_{i}}_{i}(t_{k})^{T}\textbf{R}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}\textbf{p}^{\mathcal{O}_{j}}_{j}(t_{k})$
$-\textbf{p}^{\mathcal{O}_{i}}_{ij}(t_{0})^{T}\textbf{R}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}\textbf{u}^{\mathcal{O}_{j}}_{j}(t_{k})-\textbf{p}^{\mathcal{O}_{i}}_{i}(t_{k})^{T}\textbf{R}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}\textbf{u}^{\mathcal{O}_{j}}_{j}(t_{k})$ (7)
$+\textbf{u}^{\mathcal{O}_{j}}_{j}(t_{k})^{T}\textbf{p}^{\mathcal{O}_{j}}_{j}(t_{k})=\frac{1}{2}\Big{(}d_{ij}(t_{k})^{2}-\textbf{u}^{\mathcal{O}_{i}}_{i}(t_{k})^{T}\textbf{u}^{\mathcal{O}_{i}}_{i}(t_{k})$
$-\textbf{u}^{\mathcal{O}_{j}}_{j}(t_{k})^{T}\textbf{u}^{\mathcal{O}_{j}}_{j}(t_{k})+d_{ij}(t_{k+1})^{2}\Big{)}+\textbf{u}^{\mathcal{O}_{i}}_{i}(t_{k})^{T}\textbf{R}^{\mathcal{O}_{i}}\textbf{u}^{\mathcal{O}_{j}}_{j}(t_{k})^{T}.$
Second, the obtainable terms and unknown parameters in (III-A) are organized, and one can see,
$\bar{y}_{ij}(t_{k})=$ $\textbf{u}^{\mathcal{O}_{i}}_{i}(t_{k})^{T}\textbf{p}^{\mathcal{O}_{i}}_{ij}(t_{0})-\textbf{u}^{\mathcal{O}_{j}}_{j}(t_{k})^{T}\left(\textbf{R}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}\right)^{T}\textbf{p}^{\mathcal{O}_{i}}_{ij}(t_{0})$
$-\textbf{u}^{\mathcal{O}_{i}}_{i,h}(t_{k})^{T}\textbf{R}^{\mathcal{O}_{i,h}}_{\mathcal{O}_{j,h}}\textbf{p}^{\mathcal{O}_{j}}_{j,h}(t_{k})-\textbf{p}^{\mathcal{O}_{i}}_{i,h}(t_{k})^{T}\textbf{R}^{\mathcal{O}_{i,h}}_{\mathcal{O}_{j,h}}\textbf{u}^{\mathcal{O}_{j}}_{j,h}(t_{k})$
$-\textbf{u}^{\mathcal{O}_{i}}_{i,h}(t_{k})^{T}\textbf{R}^{\mathcal{O}_{i,h}}_{\mathcal{O}_{j,h}}\textbf{u}^{\mathcal{O}_{j}}_{j,h}(t_{k}),$ (8)
where $\bar{y}_{ij}(t_{k})$ is defined as
$\bar{y}_{ij}(t_{k})\triangleq\frac{1}{2}\Big{(}d_{ij}(t_{k+1})^{2}+d_{ij}(t_{k})^{2}-\textbf{u}^{\mathcal{O}_{i}}_{i}(t_{k})^{T}\textbf{u}^{\mathcal{O}_{i}}_{i}(t_{k})$
$-\textbf{u}^{\mathcal{O}_{j}}_{j}(t_{k})^{T}\textbf{u}^{\mathcal{O}_{j}}_{j}(t_{k})\Big{)}-\textbf{u}^{\mathcal{O}_{i}}_{i}(t_{k})^{T}\textbf{p}^{\mathcal{O}_{i}}_{i}(t_{k})-\textbf{u}^{\mathcal{O}_{j}}_{j}(t_{k})^{T}\textbf{p}^{\mathcal{O}_{j}}_{j}(t_{k})$
$+u^{\mathcal{O}_{i}}_{i,z}(t_{k})p^{\mathcal{O}_{j}}_{j,z}(t_{k})+p^{\mathcal{O}_{i}}_{i,z}(t_{k})p^{\mathcal{O}_{j}}_{j,z}(t_{k})+u^{\mathcal{O}_{i}}_{i,z}(t_{k})u^{\mathcal{O}_{j}}_{j,z}(t_{k}).$
Here, we point out that $\bar{y}_{ij}(t_{k})$ is obtainable at time $t_{k+1}$. Expand the third item in the right hand of Eq. (III-A) and it has, One can check the following equations holds:
$\textbf{u}^{\mathcal{O}_{i}}_{i,h}(t_{k})^{T}\textbf{R}^{\mathcal{O}_{i,h}}_{\mathcal{O}_{j,h}}\textbf{p}^{\mathcal{O}_{j}}_{j,h}(t_{k})$ (9)
$=$ $\begin{bmatrix}\textbf{u}^{\mathcal{O}_{i}}_{i,h}(t_{k})^{T}\textbf{p}^{\mathcal{O}_{j}}_{j,h}(t_{k}),&\textbf{p}^{\mathcal{O}_{j}}_{j,h}(t_{k})\times\textbf{u}^{\mathcal{O}_{i}}_{i,h}(t_{k})\end{bmatrix}\begin{bmatrix}\cos\theta_{\mathcal{O}_{j}}^{\mathcal{O}_{i}}(t_{0})\\ \sin\theta_{\mathcal{O}_{j}}^{\mathcal{O}_{i}}(t_{0})\end{bmatrix}$
$\textbf{p}^{\mathcal{O}_{i}}_{i,h}(t_{k})^{T}\textbf{R}^{\mathcal{O}_{i,h}}_{\mathcal{O}_{j,h}}\textbf{u}^{\mathcal{O}_{j}}_{j,h}(t_{k})$
$=$ $\begin{bmatrix}\textbf{p}^{\mathcal{O}_{i}}_{i,h}(t_{k})^{T}\textbf{u}^{\mathcal{O}_{j}}_{j,h}(t_{k}),&\textbf{u}^{\mathcal{O}_{j}}_{j,h}(t_{k})\times\textbf{p}^{\mathcal{O}_{i}}_{i,h}(t_{k})\end{bmatrix}\begin{bmatrix}\cos\theta_{\mathcal{O}_{j}}^{\mathcal{O}_{i}}(t_{0})\\ \sin\theta_{\mathcal{O}_{j}}^{\mathcal{O}_{i}}(t_{0})\end{bmatrix}.$
$\textbf{u}^{\mathcal{O}_{i}}_{i,h}(t_{k})^{T}\textbf{R}^{\mathcal{O}_{i,h}}_{\mathcal{O}_{j,h}}\textbf{u}^{\mathcal{O}_{j}}_{j,h}(t_{k})$
$=$ $\begin{bmatrix}\textbf{u}^{\mathcal{O}_{i}}_{i,h}(t_{k})^{T}\textbf{u}^{\mathcal{O}_{j}}_{j,h}(t_{k}),&\textbf{u}^{\mathcal{O}_{j}}_{j,h}(t_{k})\times\textbf{u}^{\mathcal{O}_{i}}_{i,h}(t_{k})\end{bmatrix}\begin{bmatrix}\cos\theta_{\mathcal{O}_{j}}^{\mathcal{O}_{i}}(t_{0})\\ \sin\theta_{\mathcal{O}_{j}}^{\mathcal{O}_{i}}(t_{0})\end{bmatrix}.$
According to Eq. (9), define $\mathbf{\Psi}_{ij}$ and $\mathbf{\Theta}_{ij}$ as
$\mathbf{\Psi}_{ij}(t_{k})=[$ $\textbf{u}^{\mathcal{O}_{i}}_{i,h}(t_{k})^{T},u^{\mathcal{O}_{i}}_{i,z}(t_{k})-u^{\mathcal{O}_{j}}_{j,z}(t_{k}),-\textbf{u}^{\mathcal{O}_{j}}_{j,h}(t_{k})^{T},$ (10)
$-a_{ij}(t_{k}),-b_{ij}(t_{k})]^{T}$
$\mathbf{\Theta}_{ij}=[$ $\textbf{p}^{\mathcal{O}_{i}}_{ij}(t_{0})^{T},(\textbf{R}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}})^{T}(\textbf{p}^{\mathcal{O}_{i}}_{i,h}(t_{0})-\textbf{p}^{\mathcal{O}_{j}}_{j,h}(t_{0})),$
$\cos\theta_{\mathcal{O}_{j}}^{\mathcal{O}_{i}}(t_{0}),\sin\theta_{\mathcal{O}_{j}}^{\mathcal{O}_{i}}(t_{0})]^{T}.$
where $a_{ij}(t_{k})$ and $b_{ij}(t_{k})$ are defined as
$a_{ij}(t_{k})\triangleq$ $\textbf{u}^{\mathcal{O}_{i}}_{i,h}(t_{k})^{T}\textbf{p}^{\mathcal{O}_{j}}_{j,h}(t_{k})+\textbf{p}^{\mathcal{O}_{i}}_{i,h}(t_{k})^{T}\textbf{u}^{\mathcal{O}_{j}}_{j,h}(t_{k})$
$+\textbf{u}^{\mathcal{O}_{i}}_{i,h}(t_{k})^{T}\textbf{u}^{\mathcal{O}_{j}}_{j,h}(t_{k})$
$b_{ij}(t_{k})\triangleq$ $\textbf{p}^{\mathcal{O}_{j}}_{j,h}(t_{k})\times\textbf{u}^{\mathcal{O}_{i}}_{i,h}(t_{k})+\textbf{u}^{\mathcal{O}_{j}}_{j,h}(t_{k})\times\textbf{p}^{\mathcal{O}_{i}}_{i,h}(t_{k})$
$+\textbf{u}^{\mathcal{O}_{j}}_{j,h}(t_{k})\times\textbf{u}^{\mathcal{O}_{i}}_{i,h}(t_{k}).$
Note that both $a_{ij}(t_{k})$ and $b_{ij}(t_{k})$ are available at time $t_{k}$. According to Eq. (10), (III-A) can be simplified as:
$y_{ij}(t_{k})\triangleq\frac{\bar{y}_{ij}(t_{k})}{\lVert\mathbf{\Psi}_{ij}(t_{k})\rVert}=\mathbf{\Theta}_{ij}^{T}\frac{\mathbf{\Psi}_{ij}(t_{k})}{\lVert\mathbf{\Psi}_{ij}(t_{k})\rVert}\triangleq\mathbf{\Theta}_{ij}^{T}\mathbf{\Phi}_{ij}(t_{k}).$
Denote $\hat{\mathbf{\Theta}}_{ij}$ as an estimation of $\mathbf{\Theta}_{ij}$ and $\tilde{\mathbf{\Theta}}_{ij}=\hat{\mathbf{\Theta}}_{ij}-\mathbf{\Theta}_{ij}$ as the estimation error. The innovation $\epsilon_{ij}(t_{k})$ is defined as
$\epsilon_{ij}(t_{k})=\hat{\mathbf{\Theta}}_{ij}(t_{k})\mathbf{\Phi}_{ij}(t_{k})-y_{ij}(t_{k})=\tilde{\mathbf{\Theta}}_{ij}(t_{k})\mathbf{\Phi}_{ij}(t_{k}).$ (12)
The innovation $\epsilon_{ij}(t_{k})$ will be used later to design the concurrent-learning based estimator in Section III-B.
III-B Relative Localization to Neighboring Robot
Based on the innovation $\epsilon_{ij}(t_{k})$ defined in (12), now we can design the concurrent-learning based relative pose estimator to update $\hat{\mathbf{\Theta}}_{ij}(t)$ as
[…]
with the learning rate $\eta=\frac{\lambda_{min}(\textbf{S}_{ij})}{(\lambda_{max}(\textbf{U}_{ij}(t_{k}))+\lambda_{max}(\textbf{S}_{ij})^{2})}$ where matrices $\textbf{U}_{ij}(t_{k})$ and $\textbf{S}_{ij}$ are defined as $\textbf{U}_{ij}(t_{k})=\mathbf{\Phi}_{ij}(t_{k})\mathbf{\Phi}_{ij}(t_{k})^{T}\in\mathbb{R}^{7\times 7}$ and $\textbf{S}_{ij}(t_{k})=\mathbf{R}_{ij}\mathbf{R}_{ij}^{T}\in\mathbb{R}^{7\times 7}$, respectively. Here, $\mathbf{R}_{ij}=[\mathbf{\Phi}_{ij}(t_{1}),...,\mathbf{\Phi}_{ij}(t_{\varsigma})]$. For brief, define $\hat{\textbf{p}}^{\mathcal{O}_{i}}_{ij,0}\triangleq\hat{\mathbf{\Theta}}_{ij}[1:3]$ as the estimate of the initial relative position $\textbf{p}^{\mathcal{O}_{i}}_{ij}(t_{0})$ in frame $\mathcal{O}_{i}$. Similarly, the sine $\sin\theta_{\mathcal{O}_{j}}^{\mathcal{O}_{i}}(t_{0})$ and cosine $\cos\theta_{\mathcal{O}_{j}}^{\mathcal{O}_{i}}(t_{0})$ value of the initial relative frame angle can be estimated by $\hat{\mathbf{\Theta}}_{ij}[6]$ and $\hat{\mathbf{\Theta}}_{ij}[7]$. Based on estimator (13), the initial rotation matrix $\textbf{R}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}(t)$ between robot $i$ and robot $j$ can be estimated by
$\hat{\textbf{R}}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}(t):=\mathrm{Norm}\left(\begin{bmatrix}\hat{\mathbf{\Theta}}_{ij}[6],&-\hat{\mathbf{\Theta}}_{ij}[7],&0\\ \hat{\mathbf{\Theta}}_{ij}[7],&\hat{\mathbf{\Theta}}_{ij}[6],&0\\ 0,&0,&1\end{bmatrix}\right),$ (14)
where $R=\mathrm{Norm}(A)$ is a function that normalizes an antisymmetric matrix $A$ into a rotation matrix $R$.
It is worth mentioning that for 2D robots, there is no need to estimate the relative position in $z$-axis, the variables in $z$ dimension in Eq. (10) can be set to zero. The estimator for the two-dimensional situation can be easily obtained based on Eq. (13), and we omit it for brief. In the next theorem, we will show that the exponential convergence of the proposed estimator (13) can be guaranteed if the recorded data matrix $S_{ij}$ satisfies the following assumption.
Assumption 1.
The recorded data matrix $S_{ij}$ is full rank.
Assumption 1 indicates that the data matrix $S_{ij}$ records sufficient information, which is also known as SE condition. Compared to the PE condition, SE condition is much weaker and can also be found in many data-driven estimation results like [30] and [31]. We also analyze the observability of relative pose when the relative motion between neighboring robots is different, which can be found in Appendix 15-D. The main results regarding the convergence of the estimator (13) are presented in the following theorem.
Theorem 1.
Under Assumption 1, estimation error $\tilde{\mathbf{\Theta}}_{ij}$ is globally exponentially stable with relative pose estimator (13).
Proof.
See Appendix 15-A ∎
Remark 1.
Different from the latest result [1] in which the relative position should be available to estimate the unknown relative orientation, the estimator (13) proposed in this article can achieve inter-robot relative pose with only UWB distance and IO displacement measurements. Theorem 1 indicates that when the historical measurement data records sufficient information, i.e. data matrix $S_{ij}$ has full rank, the localization error converges exponentially.
In application, both UWB and odometer sensors have measurement errors. With the effective outlier detection algorithms, most measurement errors can be considered that the measurement error is bounded. According to Theorem 1 in [32], when the measurement error is bounded, there is a linear relationship between the estimation error and the upper bound of the measurement error.
III-C Relative Localization to Leader Robot
To achieve the formation control, the relative pose between each robot $i$ and the leader robot are required. However, due to the directed communication topology, not all the robots has the ability to estimate the relative pose to the leader robot directly. To deal with this issue, each robot need to utilize the information of its neighboring robots to estimate its own relative pose to the leader robot. To achieve the cooperative pose estimation task, we assume that the communication topology graph $\mathcal{G}$ is directed acyclic graph (DAG), such as the graph shown in Fig. 2 B. This condition is mild and can be found in many formation control results, such as [11] and [33]. For large-scale robot swarms, the hierarchical layering structure of swarm has been proven to be an effective approach to achieve cooperative control task in the latest result [34].
Figure 2: An example directed acyclic graph, the leader robot and follower robots are denoted in the orange and blue dots.
In our approach, we employ a DAG hierarchical topology to avoid robots using neighbor information that is too distant from the leader within the topology layer, as this can introduce additional localization errors. For instance, consider the measurement topology depicted in Fig. 2A. The measurement topology can be reorganized into the DAG form shown in Fig. 2B through a hop based distributed neighbor selection algorithm[35]. In the DAG graph, Robot 2 directly locates Leader 0 without relying on the estimation information of robot 3 and robot 5, thereby avoiding the introduction of additional localization errors. Moreover, with the distributed neighbor selection algorithm[35], the localization task can still be achieved even if some edges fail, as long as the entire topology remains directionally connected.
The estimators $\hat{\textbf{q}}^{\mathcal{O}_{i}}_{i0}$ and $\hat{\textbf{Q}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}$ are designed to estimate the initial relative position and frame rotation matrix between robot $i$ and the leader robot. Define the estimation error as $\tilde{\textbf{q}}^{\mathcal{O}_{i}}_{i0}=\hat{\textbf{q}}^{\mathcal{O}_{i}}_{i0}-\textbf{p}^{\mathcal{O}_{i}}_{i0}$ and $\tilde{\textbf{Q}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}=\hat{\textbf{Q}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}-\textbf{R}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}$. For robot $i$ in Layer $1$, i.e. $0\in\mathcal{N}_{i}$, $\hat{\textbf{q}}^{\mathcal{O}_{i}}_{i0,0}$ and $\hat{\textbf{Q}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}$ are designed as,
$\hat{\textbf{q}}^{\mathcal{O}_{i}}_{i0,0}(t)=\hat{\textbf{p}}^{\mathcal{O}_{i}}_{i0,0}(t),\quad\hat{\textbf{Q}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}(t)=\hat{\textbf{R}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}(t).$ (15)
For robot $i$ in other Layers, i.e. $0\notin\mathcal{N}_{i}$, $\hat{\textbf{q}}^{\mathcal{O}_{i}}_{i0,0}$ and $\hat{\textbf{Q}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}$ are designed as,
$\hat{\textbf{q}}^{\mathcal{O}_{i}}_{i0,0}(t)$ $=\sum_{j\in\mathcal{N}_{i}}\frac{1}{|\mathcal{N}_{i}|}\left(\hat{\textbf{p}}^{\mathcal{O}_{i}}_{ij,0}(t)+\hat{\textbf{R}}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}(t)\hat{\textbf{q}}^{\mathcal{O}_{j}}_{j0,0}(t)\right)$ (16)
$\hat{\textbf{Q}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}(t)$ $=\mathrm{Norm}\left(\sum_{j\in\mathcal{N}_{i}}\frac{1}{|\mathcal{N}_{i}|}\left(\hat{\textbf{R}}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}(t)\hat{\textbf{Q}}^{\mathcal{O}_{j}}_{\mathcal{O}_{0}}(t)\right)\right).$
$\triangleq\begin{bmatrix}\widehat{\cos\theta_{\mathcal{O}_{0}}^{\mathcal{O}_{i}}}(t),&-\widehat{\sin\theta_{\mathcal{O}_{0}}^{\mathcal{O}_{i}}}(t),&0\\ \widehat{\sin\theta_{\mathcal{O}_{0}}^{\mathcal{O}_{i}}}(t),&\widehat{\cos\theta_{\mathcal{O}_{0}}^{\mathcal{O}_{i}}}(t),&0\\ 0,&0,&1\end{bmatrix}.$
With estimators $\hat{\textbf{q}}^{\mathcal{O}_{i}}_{i0,0}(t)$ and $\hat{\textbf{Q}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}(t)$ in (15) and (16), relative pose estimators $\hat{\textbf{q}}^{\mathcal{O}_{i}}_{i0}(t)$ and $\hat{\textbf{Q}}^{\Sigma_{i}}_{\Sigma_{0}}(t)$ are defined as
$\hat{\textbf{q}}^{\mathcal{O}_{i}}_{i0}(t)\triangleq\hat{\textbf{q}}^{\mathcal{O}_{i}}_{i0,0}(t)+\textbf{z}^{\mathcal{O}_{i}}_{i}(t)-\hat{\textbf{Q}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}\textbf{z}^{\mathcal{O}_{0}}_{0}(t)$ (17)
$\hat{\textbf{Q}}^{\Sigma_{0}}_{\Sigma_{i}}(t)\triangleq\begin{bmatrix}\widehat{\cos\theta_{\Sigma_{i}}^{\Sigma_{0}}}(t),&-\widehat{\sin\theta_{\Sigma_{i}}^{\Sigma_{0}}}(t),&0\\ \widehat{\sin\theta_{\Sigma_{i}}^{\Sigma_{0}}}(t),&\widehat{\cos\theta_{\Sigma_{i}}^{\Sigma_{0}}}(t),&0\\ 0,&0,&1\end{bmatrix},$
where $\widehat{\cos\theta_{\Sigma_{i}}^{\Sigma_{0}}}(t)$ and $\widehat{\sin\theta_{\Sigma_{i}}^{\Sigma_{0}}}(t)$ are defined as
$\widehat{\cos\theta_{\Sigma_{i}}^{\Sigma_{0}}}(t)\triangleq$ $\widehat{\cos\theta_{\mathcal{O}_{0}}^{\mathcal{O}_{i}}}(t)\cos(\phi_{\Sigma_{i}}^{\mathcal{O}_{i}}(t)-\phi_{\Sigma_{0}}^{\mathcal{O}_{0}}(t))$
$+\widehat{\sin\theta_{\mathcal{O}_{0}}^{\mathcal{O}_{i}}}(t)\sin(\phi_{\Sigma_{i}}^{\mathcal{O}_{i}(t)}-\phi_{\Sigma_{0}}^{\mathcal{O}_{0}}(t))$
$\widehat{\sin\theta_{\Sigma_{i}}^{\Sigma_{0}}}(t)\triangleq$ $\widehat{\cos\theta_{\mathcal{O}_{0}}^{\mathcal{O}_{i}}}(t)\sin(\phi_{\Sigma_{i}}^{\mathcal{O}_{i}(t)}-\phi_{\Sigma_{0}}^{\mathcal{O}_{0}}(t))$
$-\widehat{\sin\theta_{\mathcal{O}_{0}}^{\mathcal{O}_{i}}}(t)\cos(\phi_{\Sigma_{i}}^{\mathcal{O}_{i}(t)}-\phi_{\Sigma_{0}}^{\mathcal{O}_{0}}(t)).$
We point out that the odometry information $\textbf{z}^{\mathcal{O}_{i}}_{i}(t)$ of leader in (17) can be obtained by integrating $v_{0,h}$, $v_{0,z}$ and $w_{0}$ with the following Assumption 2. Besides, $\mathbf{z}^{\mathcal{O}_{i}}_{i}(t)$ can also be transmitted between robots via communication [1].
Assumption 2.
The velocity $v_{0,h}$, $v_{0,v}$ and angle velocity $w_{0}$ of the leader robot are bounded and available to all followers.
Note that Assumption 2 indicates that each robot has some prior knowledge of the leader’s movements, which is reasonable and can also be found in cooperation localization works [19, 1]. The convergence of estimator (15) and (16) are presented in the following theorem.
Theorem 2.
Under Assumption 2, given $(i,j)\in\mathcal{G}$, the record data matrix $S_{ij}$ satisfies Assumption 1. Then for robot $i$ in each layer, the cooperative localization errors $\tilde{\textbf{q}}^{\mathcal{O}_{i}}_{i0,0}$ and $\tilde{\textbf{Q}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}$ converge asymptotically.
Proof.
See Appendix 15-B ∎
III-D Distributed Formation tracking Controller
The real-time relative orientation for robot $i$ in the leader’s odometer frame $\mathcal{O}_{0}$ can be defined as
$\theta_{\Sigma_{i}}^{\mathcal{O}_{0}}(t)=\phi_{\Sigma_{i}}^{\mathcal{O}_{i}}(t)-\theta_{\mathcal{O}_{0}}^{\mathcal{O}_{i}}(t_{0}),$ (19)
where $\theta_{\mathcal{O}_{0}}^{\mathcal{O}_{i}}(t_{0})$ is the initial relative orientation to the leader robot. Define the formation error $\textbf{e}_{i}=[\textbf{e}_{i,p}^{T},e_{i,c},e_{i,s}]^{T}$ as
$\textbf{e}_{i,p}$ $=\textbf{R}_{\mathcal{O}_{0}}^{\Sigma_{i}}\left(\textbf{p}^{\mathcal{O}_{0}}_{i0}-\textbf{p}_{i}^{*}\right)$ (20)
$e_{i,c}$ $=1-\cos\theta_{\Sigma_{i}}^{\Sigma_{0}},\quad e_{i,s}=\sin\theta_{\Sigma_{i}}^{\Sigma_{0}},$
where $\textbf{e}_{i,p}\triangleq[e_{i,x},e_{i,y},e_{i,z}]^{T}$. We note that although the value of $\textbf{p}^{\mathcal{O}_{0}}_{i0}$, $\sin\theta_{\mathcal{O}_{0}}^{\mathcal{O}_{i}}(t_{0})$ and $\cos\theta_{\mathcal{O}_{0}}^{\mathcal{O}_{i}}(t_{0})$ in (10) can not be measured directly, they can be estimated asymptotically with the estimator (17). Define the estimated formation tracking error for robot $i$ as
$\hat{\textbf{e}}_{i,p}$ $=\hat{\textbf{R}}_{\mathcal{O}_{0}}^{\Sigma_{i}}\left(\left(\hat{\textbf{Q}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}(t)\right)^{T}\hat{\textbf{q}}^{\mathcal{O}_{i}}_{i0}(t)-\textbf{p}_{i}^{*}\right)$
$\hat{e}_{i,c}$ $=1-\widehat{\cos\theta_{\Sigma_{i}}^{\Sigma_{0}}}(t),\quad\hat{e}_{i,s}=\widehat{\sin\theta_{\Sigma_{i}}^{\Sigma_{0}}}(t)$ (21)
$\hat{\textbf{R}}_{\mathcal{O}_{0}}^{\Sigma_{i}}$ $\triangleq\begin{bmatrix}\widehat{\cos\theta_{\Sigma_{i}}^{\mathcal{O}_{0}}}(t),&\widehat{\sin\theta_{\Sigma_{i}}^{\mathcal{O}_{0}}}(t),&0\\ -\widehat{\sin\theta_{\Sigma_{i}}^{\mathcal{O}_{0}}}(t),&\widehat{\cos\theta_{\Sigma_{i}}^{\mathcal{O}_{0}}}(t),&0\\ 0,&0,&1\end{bmatrix},$
where $\widehat{\cos\theta_{\Sigma_{i}}^{\mathcal{O}_{0}}}(t)$ and $\widehat{\sin\theta_{\Sigma_{i}}^{\mathcal{O}_{0}}}(t)$ are defined as
$\widehat{\cos\theta_{\Sigma_{i}}^{\mathcal{O}_{0}}}(t)$ $=\widehat{\cos\theta_{\mathcal{O}_{0}}^{\mathcal{O}_{i}}}(t)\cos\phi_{\Sigma_{i}}^{\mathcal{O}_{i}}(t)+\widehat{\sin\theta_{\mathcal{O}_{0}}^{\mathcal{O}_{i}}}(t)\sin\phi_{\Sigma_{i}}^{\mathcal{O}_{i}}(t)$
$\widehat{\sin\theta_{\Sigma_{i}}^{\mathcal{O}_{0}}}(t)$ $=\widehat{\cos\theta_{\mathcal{O}_{0}}^{\mathcal{O}_{i}}}(t)\sin\phi_{\Sigma_{i}}^{\mathcal{O}_{i}}(t)-\widehat{\sin\theta_{\mathcal{O}_{0}}^{\mathcal{O}_{i}}}(t)\cos\phi_{\Sigma_{i}}^{\mathcal{O}_{i}}(t).$
[…]
$\tilde{\textbf{e}}_{i}=[\tilde{e}_{i,x},\tilde{e}_{i,y},\tilde{e}_{i,c},\tilde{e}_{i,s}].$ (22)
The proposed control scheme consists of two stage including localization stage and sharp formation stage. Initially, each robot $i$ performs specified circular actions [36] to achieve localization enhancement in the localization stage, i.e. collects measurement data matrix $\textbf{S}_{ij}$ with its neighbor $j$. In this stage, for the follower robot $i$, the control input is designed as
[…]
where $r_{i}$ is the circular radius, $c_{i,v}$ and $c_{i,w}$are positive constant. For robot $i$, the first control stage ends when all $S_{ij},j\in\mathbb{N}_{i}$ data matrices are full rank. The second control stage begin when all the robots end their first stages. When the second stage control start can be determined with methods such as hop-count propagation [37]. In the second control stage, for the follower robot $i$, the control input is designed as
[…]
where $k1>0$, $k2>0$, $k3>0$ and $k4>0$ are the control gain. Note that only the estimated formation tracking errors (III-D) are adopoted in controller (24). To achieve the formation control task, the following assumption should be satisfied,
Assumption 3.
The angle velocity $w_{0}$ is persistently exciting for all $t\geq t_{0}$.
Figure 3: Pose estimation error defined in (22) under different noise level.
Figure 4: Convergence time (time required for the relative error of the estimated value to be less than 5%) of the two algorithm under different swarm scales.
Figure 5: Smoothness indicator defined in Eq.(25) under different swarm scale.
It is worth to point out that the persistently exciting angle velocity $w_{0}$ is a common condition to achieve asymptotically tracking control for nonholonomic robots and can be found in related results [38][1]. Furthermore, the requirement for $w_{0}$ to meet the PE condition is only for the asymptotically tracking control of nonholonomic robot systems, whereas the convergence of the proposed estimators (13) or (17) does not require this condition.
To better illustrate the calculation process of algorithm data collection, collaborative localization, and control instructions, we have added pseudocode for the proposed method in Appendix A-B. The stability of the proposed controller can be given in the following theorem.
Theorem 3.
Under Assumption 2 and 3, given $(i,j)\in\mathcal{G}$, the recorded data matrix $S_{ij}$ satisfies Assumption 1. consider the closed-loop system consisting of $N+1$ nonholonomic robots (1), the relative pose estimator (17) and local controller (24), then the tracking error $\textbf{e}_{i}$ asymptotically converge to zero.
Proof.
See Appendix 15-C ∎
Figure 6: Estimation accuracy of localization method under different sensor outlier probabilities.
Figure 7: Detection success rate of outlier detection algorithm under different sensor outlier probabilities.
IV Simulation Results
In this section, we present the simulation results to demonstrate the effectiveness and robustness of the proposed method.
IV-A Performance of the PE-free relative localization method
Firstly, we evaluated the noisy robustness of the proposed method and the latest method [14]. Under the two-robot relative localization case, we add standard deviations ranging from 0.01m to 1m to each measurement of distance and odometer displacement. For each noise case, we set up 100 sets of simulation scenarios with different initial conditions. The parameters used in the simulation are $\Delta t=0.05$, $r_{0}=2$, $r_{0}=3$, $c_{0,w}=0.1$, $c_{1,w}=0.3$, $c_{0,v}=0.2$, $c_{1,v}=0.3$, $k_{1}=1$, $k_{2}=0.5$, $k_{3}=0.4$, $k_{4}=0.2$. For more information about parameter selection, please refer to Appendix B-A. Parameters adopted in the contrast simulation can be found in [14]. The simulation results are shown in Fig. 3. It can be seen that as the measurement noise level increases, the localization accuracy of both algorithms decreases. However, the performance of the proposed data-driven algorithm is still inferior to that of the PE condition based method in [14]. The fundamental reason is that the data-driven method can tolerate abnormal measurement values to a certain extent.
Secondly, we evaluated the convergence speed and control performance of the proposed method and the latest method [14]. Under different swarm scales, we set up 100 sets of formation simulation scenarios with different initial conditions. The parameters used in the simulation are $\Delta t=0.05$, $r_{i}=2+2\mathrm{rand}(-0.5,0.5)$, $c_{i,w}=4\mathrm{rand}(-0.5,0.5)$, $\mathrm{rand}(0,1)$, $k_{1}=1$, $k_{2}=0.5$, $k_{3}=0.4$, $k_{4}=0.2$. Note that function $\mathrm{rand}(a,b)$ represents the random value from $a$ to $b$. In the comparative simulation, except for the localization and control algorithms, all other conditions were completely the same. The control performance is reflected by the smoothness of the motion trajectory. Define the smoothness indicator as:
$S_{i}=\int_{t_{0}}^{t_{\mathrm{end}}}\lVert v_{i}-v_{\mathrm{leader}}\rVert$ (25)
where $t_{0}$ and $t_{\mathrm{end}}$ denote the beginning time and ending time of the simulation, $v_{i}$ is the velocity of robot $i$, $v_{\mathrm{leader}}$ is the velocity of the leader robot. The smoothness indicator(25) represents the degree to which the speed of robot $i$ deviates from the expected speed (i.e. leader speed). According to this definition, the convergence time and smoothness indicator(25) of the two algorithms is shown in Fig. 4 and Fig. 5. It can be seen that our algorithm generally has better convergence speed than existing results in swarms of different sizes. This is because the data-driven localization algorithm adopted both historical information and current information, which accelerates the convergence speed. Furthermore, one can see that the proposed method outperforms existing method [14] based on PE condition in terms of trajectory smoothness across various swarm sizes. This is because the PE condition based localization method [14] requires the relative displacement between robots satisfy the PE condition and add noise and other PE signals to the speed command. When the swarm size is 5, in a simulation experiment, the robot motion trajectories of two algorithms are shown in Appendix Appendix B-B. It can be clearly seen that our algorithm outperforms the PE based algorithm in terms of control performance (trajectory smoothness).
IV-B Outlier measurement value detection
Figure 8: Diagram of the experimental system.
Figure 9: Snapshots of the grounded robot formation control.
Figure 10: Tracking error of each robot. A. Position tracking error $e_{x,i}$ and $e_{y,i}$ defined in (10) of each robot. B. Orientation tracking error $e_{c,i}$ and $e_{s,i}$ defined in (10) of each robot.
To reduce the impact of measurement outliers, an outlier detection method is proposed in Appendix A-B. The core idea of the detection method is that for any nominal distance measurement $d_{ij}[k]$, the absolute value of the difference between $d_{ij}[k]$ and any historical distance $d_{ij}[m]$ should be less than the sum of the robot’s displacement between the two measurements, i.e. $\lVert z_{i}^{\mathcal{O}_{i}}[k]-z_{i}^{\mathcal{O}_{i}}[m]\rVert+\lVert z_{j}^{\mathcal{O}_{j}}[k]-z_{j}^{\mathcal{O}_{j}}[m]\rVert$, which is actually a triangular inequality. 100 sets of relative localization experiments under different initial conditions are conducted to verify the effectiveness of the outlier detection trick. To simulate the actual situation, we added measurement noise with a standard deviation of 0.05 to the UWB and IO measurements. The UWB sensor has a certain probability of detecting outlier values each time, that is, measurement noise with a standard deviation of 3. The probabilities of outlier values occurring are 0.05, 0.1, 0.2, 0.4, 0.6, and 0.8, respectively. The parameters adopted are the same as those in the two-robot case in Section IV-A. The pose estimation error with and without outlier detection algorithms is shown in Fig. 6. The accuracy of outlier detection is shown in Fig. 7. From the simulation results, one can see that the outlier detection algorithms can effectively detect measurement outliers and improve the accuracy of localization algorithms.
Due to space limitations, we provide more simulation results in Appendix B-(C-E) of the extended version of this article, including validation of the effectiveness of the algorithm with a large scale of swarm, noise robustness, and sensor failures.
V Experiment results
V-A Experiment Setup
In Fig. 8, the diagram of the experimental system is illustrated. The platform for 3D experiments is our unmanned aerial vehicle built based on a PX4 flight controller. A kind of nonholonomic robot Turtlebot3 is adopted as the experimental platform for the 2D experiments. Each robot is equipped with an IO to measure its own displacement and a UWB sensor to measure the real-time distance to its neighbors. For the drone, a Visual inertial odometer (VIO) is adopted to improve the accuracy of IO. Note that VIO is a technically mature odometer that is also used in related distance and odometer based relative localization work [12, 10]. There are several threads running on a ground working station, each thread $i$ receives the real-time measurement information sent by robot $i$. Distributed communication between threads are imitated with robot operating system (ROS), the communication topology is also defined in Fig. 8. The proposed distributed relative localization and distributed formation controller for robot $i$ are also executed by thread $i$. It should be noted that despite the usage of the ground station in the experiments, the algorithm still runs in a distributed manner.
Figure 11: Estimation error of each robot. A. Position estimate error $\tilde{e}_{x,i}$ and $\tilde{e}_{y,i}$ defined in (22) of each robot. B. Orientation estimate error $\tilde{e}_{c,i}$ and $\tilde{e}_{s,i}$ defined in (22) of each robot.
Figure 12: Pose estimation and tracking error in ten experiments with different initial position conditions.
Figure 13: Snapshots of the aerial robot relative tracking experiments.
Figure 14: Estimation error and robot tracking error of robot $1$. A. Estimated error $\tilde{\textbf{e}}_{1}$ defined in (22) of robot $1$. B. Tracking error $\textbf{e}_{1}$ defined in (10) of robot $1$.
V-B Grounded robot Formation Control Experiment
Multi-robot formation results are provided to verify the effectiveness of the localization and control method in 2D situation. In this experiment, $3$ follower robots and a leader robot are assigned to achieve the formation task with the prescribed relative position $\textbf{p}_{1}^{*}=[0.5,-0.5,0]^{T}$, $\textbf{p}_{2}^{*}=[-0.5,-0.5,0]^{T}$ and $\textbf{p}_{3}^{*}=[0,0.5,0]^{T}$ in leader odometer frame $\mathcal{O}_{i}$. Parameter used in the experiment are $\Delta t=0.1$, $r_{i}=0.3,i=0,...,3$, $c_{0,w}=0.5$, $c_{1,w}=0.2$, $c_{2,w}=0.2$, $c_{3,w}=0.5$, $k_{1}=1$, $k_{2}=1$, $k_{3}=0.5$, $k_{4}=0$. After the data collection stage, the leader robot maintain a constant velocity $v_{0}$ and angle velocity $w_{0}$ to achieve a circular formation.
In this experiment, NOKOV motion capture system is used to record the ground truth pose of the leader and follower robots. The snapshuts of the experiment process is shown in Fig. 9. The estimated error $\tilde{\textbf{e}}_{i}$ defined in (22) are shown in Fig. 11. Fig. 11 A shows the position estimate error $\tilde{e}_{x,i}$ and $\tilde{e}_{y,i}$, Fig. 11 B shows the orientation estimate error $\tilde{e}_{c,i}$ and $\tilde{e}_{s,i}$. The tracking error $\textbf{e}_{i}$ defined in (10) are shown in Fig. 10. Fig. 11 A shows the position tracking error $e_{x,i}$ and $e_{y,i}$, Fig. 11 B shows the orientation estimation error $e_{c,i}$ and $e_{s,i}$. One can see that when the measurement data is collected sufficiently, the estimation error and tracking error fast converge. The formation control task is achieved. This experiment verifies the effectiveness of the proposed localization algorithm.
To verify the consistency of the algorithm’s results under different initial conditions, we conducted Tb3 swarm formation experiments under 10 different initial conditions, with each experiment lasting 150 seconds. We recorded the average estimation and control errors throughout the experiment, as well as the relative pose estimation and tracking control error results, as shown in the Fig. 12. From the results, it can be seen that the experimental results at different initial positions have good consistency, and the differences in the results are mainly due to the possible variations in the measurement noise of the sensors in each experiment. The snapshot of experiments starting from three different initial conditions in ten sets of experiments can be found in Appendix B-F.
V-C Aerial robot Tracking Control Experiment
Both 2-D and 3-D localization and formation control experiments are provided to verify the effectiveness of the method.
In this experiment, follower robot $1$ is assigned to track the leader robot with a prescribed relative position $\textbf{p}_{1}^{*}\triangleq[1.2,1.2,0]^{T}$ in leader odometer frame $\mathcal{O}_{i}$. Parameter used in the experiment are $\Delta t=0.1$, $r_{0}=0.3$, $r_{1}=0.5$, $c_{0,w}=-0.5$, $c_{1,w}=0.4$, $k_{1}=0.4$, $k_{2}=0.8$, $k_{3}=0.5$, $k_{4}=0.5$. The condition for the end of the data collection stage is that the ratio of the minimum to maximum eigenvalues of the data collection matrix $S_{ij}$ is greater than the preset value of 0.1, which can usually be estimated through simulation.
Dual aerial robots tracking experiment results are provided to verify the effectiveness of the relative localization method in 3D situation. In this experiment, NOKOV motion capture system is used to record the ground truth pose of the leader and follower robots. The snapshots of the experiment process is shown in Fig. 13, the trajectory of the leader robot and the follower robot are described with orange and yellow curves in Fig. 13. The estimated error $\tilde{\textbf{e}}_{1}$ defined in (22) and tracking error $\textbf{e}_{1}$ defined in (10) are shown in Fig. 14 A and Fig. 14 B, respectively. One can see that when the measurement data is collected sufficiently, the estimation error and tracking error converge fast. This experiment verifies the effectiveness of the proposed localization algorithm in 3D case.
V-D Outdoor Formation Control Experiment
To demonstrate the effectiveness of our algorithm in other environments (such as the corridor you mentioned), we tested the relative localization and formation control algorithm in a building courtyard corridor environment. The parameter adopted is the same as the indoor experiment. A snapshot of the experimental process is shown in Fig. 15. From the snapshot, one can see that the robot swarm implements relative localization based formation control task based on UWB and IO measurements, verifying the effectiveness of our method.
VI Conclusion
This article has addressed the relative localization based distributed formation control for multi nonholonomic robots. The motivation arises from applications such as object transport in indoor and underground environments. First, we proposed a concurrent-learning based relative pose estimator, enabling inter-robot relative localization within local frame through only UWB distance and IO measurements. Second, to determine the relative pose between each robot and the leader robot, a cooperative localization strategy is introduced. Third, we devised the distributed formation controller based on the relative pose estimators designed in this article. With the proposed methods, distributed formation control for nonholonomic robots is achieved with only UWB distance and IO measurements. Both theoretical analysis and indoor experiments involving four robots were conducted to validate the effectiveness and adaptability of the proposed methods.
Figure 15: Snapshots of the localization and formation control in the courtyard corridor environment.
-A Proof of Theorem 1
The dynamics of the relative localization error is
$\tilde{\mathbf{\Theta}}_{ij}(t_{k+1})=\left(I-\eta U_{ij}(t_{k})-\eta S_{ij}\right)\tilde{\mathbf{\Theta}}_{ij}(t_{k}).$ (26)
Choose a Lyapunov candidate as
$V_{ij}(t_{k})=\tilde{\mathbf{\Theta}}_{ij}(t_{k})^{T}\tilde{\mathbf{\Theta}}_{ij}(t_{k}).$ (27)
Define $\Delta V_{ij}=V_{ij}(t_{k+1})-V_{ij}(t_{k})$, substituting (26) into (27), it has
$\Delta V_{ij}=$ $\tilde{\mathbf{\Theta}}_{ij}(t_{k})^{T}\left(-2\eta S_{ij}+2\eta^{2}S_{ij}U_{ij}(t_{k})+\eta^{2}S_{ij}^{2}\right)\tilde{\mathbf{\Theta}}_{ij}(t_{k})$
$-\tilde{\mathbf{\Theta}}_{ij}(t_{k})^{T}\eta U_{ij}(t_{k})(2I-\eta U_{ij}(t_{k}))\tilde{\mathbf{\Theta}}_{ij}(t_{k}).$
It can be checked that
$\Delta V_{ij}\leq$ $\left(-2\eta\lambda_{\min}(S_{ij})+2\eta^{2}\lambda_{\max}(U_{ij}(t_{k}))\lambda_{\max}(S_{ij})\right.$
$\left.+\eta^{2}\lambda_{\max}(S_{ij})^{2}\right)\lVert\tilde{\mathbf{\Theta}}_{ij}(t_{k})\rVert^{2}$
$+\eta^{2}\lambda_{\max}(U_{ij}(t_{k}))^{2}\lVert\tilde{\mathbf{\Theta}}_{ij}(t_{k})\rVert^{2}.$
With the learning rate $\eta=\frac{\lambda_{min}(\textbf{S}_{ij})}{(\lambda_{max}(\textbf{U}_{ij}(t_{k}))+\lambda_{max}(\textbf{S}_{ij})^{2})}$, one can get,
$\Delta V_{ij}\leq$ $\frac{-\lambda_{\min}(S_{ij})^{2}}{(\lambda_{\max}(U_{ij}(t_{k}))(t_{k})+\lambda_{\max}(S_{ij}))^{2}}\lVert\tilde{\mathbf{\Theta}}_{ij}(t_{k})\rVert^{2}.$ (28)
One can see that for matrix $U_{ij}(t_{k})=u_{ij}(t_{k})u_{ij}(t_{k})^{T}\in\mathbb{R}^{3}$. According to Assumption 1, it can be check that the maximum eigenvalue satisfies
$\lambda_{\max}(U_{ij}(t_{k}))=\mathbf{\Phi}_{ij}(t_{k})^{T}\mathbf{\Phi}_{ij}(t_{k})\leq 1.$ (29)
Substituting (28) into (29), the relative localization error $\tilde{\mathbf{\Theta}}_{ij}$ satisfies
$\lVert\tilde{\mathbf{\Theta}}_{ij}(t_{k})\rVert\leq\left(\sqrt{\left(1-\frac{\lambda_{\min}(S_{ij})^{2}}{1+\lambda_{\max}(S_{ij}))^{2}}\right)}\right)^{k}\lVert\tilde{\mathbf{\Theta}}_{ij}(t_{0})\rVert.$
which means that $\tilde{\mathbf{\Theta}}_{ij}$ is global exponentially stable.
-B Proof of Theorem 2
Under the DAG topology, Each robot $i$ in layer $l+1$ has at least one neighbor $j$ in layer $l$. For the robots in layer $1$, the leader state estimators $\hat{\textbf{q}}^{\mathcal{O}_{i}}_{i0,0}$ and $\hat{\textbf{Q}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}$ are directly set as the relative pose estimators $\hat{\textbf{p}}^{\mathcal{O}_{i}}_{i0,0}$ and $\hat{\textbf{R}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}$. According to Theorem 1, the estimation error $\lVert\tilde{\textbf{q}}^{\mathcal{O}_{i}}_{i0,0}\rVert$ and $\lVert\tilde{\textbf{Q}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}\rVert$ converge to zero asymptotically. Next, we will use mathematical induction to prove the convergence of $\lVert\tilde{\textbf{q}}^{\mathcal{O}_{i}}_{i0,0}\rVert$ and $\lVert\tilde{\textbf{Q}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}\rVert$. Assuming that for robot $i$ in layer $l$, the estimation error $\lVert\tilde{\textbf{q}}^{\mathcal{O}_{i}}_{i0,0}\rVert$ converges to zero asymptotically. For robot $i$ in layer $l+1$, it has
$\tilde{\textbf{q}}^{\mathcal{O}_{i}}_{i0,0}(t)=$ $\sum_{j\in\mathcal{N}_{i}}\frac{1}{|\mathcal{N}_{i}|}\left(\tilde{\textbf{p}}^{\mathcal{O}_{i}}_{ij,0}(t)-\tilde{\textbf{R}}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}(t)\textbf{p}^{\mathcal{O}_{j}}_{j0}-\hat{\textbf{R}}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}(t)\tilde{\textbf{q}}^{\mathcal{O}_{j}}_{j0}(t)\right)$
$\lVert\tilde{\textbf{q}}^{\mathcal{O}_{i}}_{i0,0}(t)\rVert\leq$ $\sum_{j\in\mathcal{N}_{i}}\frac{1}{|\mathcal{N}_{i}|}\left(\lVert\tilde{\textbf{p}}^{\mathcal{O}_{i}}_{ij,0}(t)\rVert+\lVert\tilde{\textbf{R}}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}(t)\rVert\lVert\textbf{p}^{\mathcal{O}_{j}}_{j0}\rVert\right.$
$\left.+\lVert\hat{\textbf{R}}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}(t)\rVert\lVert\tilde{\textbf{q}}^{\mathcal{O}_{j}}_{j0}(t)\rVert\right).$ (31)
One can see that the norm of the normalized rotation matrix $\lVert\hat{\textbf{R}}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}(t)\rVert$ in (-B) is bounded. Under Assumption 1 and Theorem 1, $\lVert\tilde{\textbf{p}}^{\mathcal{O}_{i}}_{ij,0}(t)\rVert$ converges to zero asymptotically. Under the assuming convergence of $\lVert\tilde{\textbf{q}}^{\mathcal{O}_{j}}_{j0,0}\rVert$ for $j$ in layer $l$, $\lVert\tilde{\textbf{q}}^{\mathcal{O}_{i}}_{i0,0}(t)\rVert$ converges asymptotically. Similarly, assuming that for robot $i$ in layer $l$, the estimation $\lVert\tilde{\textbf{Q}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}\rVert$ converges to zero asymptotically. Then for robot $i$ in layer $l+1$, it has
$\tilde{\textbf{Q}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}(t)=$ $\sum_{j\in\mathcal{N}_{i}}\frac{1}{|\mathcal{N}_{i}|}\left(\textbf{R}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}-\hat{\textbf{R}}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}(t)\hat{\textbf{Q}}^{\mathcal{O}_{j}}_{\mathcal{O}_{0}}(t)\right)$ (32)
$=$ $-\sum_{j\in\mathcal{N}_{i}}\frac{1}{|\mathcal{N}_{i}|}\left(\tilde{\textbf{R}}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}(t)\textbf{R}^{\mathcal{O}_{j}}_{\mathcal{O}_{0}}(t)+\hat{\textbf{R}}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}(t)\tilde{\textbf{Q}}^{\mathcal{O}_{j}}_{\mathcal{O}_{0}}(t)\right),$
$\lVert\tilde{\textbf{Q}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}(t)\rVert\leq$ $\sum_{j\in\mathcal{N}_{i}}\frac{1}{|\mathcal{N}_{i}|}\Big{(}\lVert\tilde{\textbf{R}}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}(t)\rVert\lVert\textbf{R}^{\mathcal{O}_{j}}_{\mathcal{O}_{0}}(t)\rVert$
$+\lVert\hat{\textbf{R}}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}(t)\rVert\lVert\tilde{\textbf{Q}}^{\mathcal{O}_{j}}_{\mathcal{O}_{0}}(t)\rVert\Big{)}.$ (33)
Under Assumption 1 and Theorem 1, $\lVert\tilde{\textbf{R}}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}(t)\rVert$ in (32) converges asymptotically. Under the assuming convergence of $\lVert\tilde{\textbf{Q}}^{\mathcal{O}_{j}}_{\mathcal{O}_{0}}(t)\rVert$ for robot $j$ in layer $l$, $\lVert\tilde{\textbf{Q}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}(t)\rVert$ converges asymptotically for robot $i$ in layer $l+1$.
-C Proof of Theorem 3
The derivative of the tracking error $\textbf{e}_{i}$ is
$\dot{\textbf{e}}_{i}=\begin{bmatrix}-v_{0,h}+v_{0,h}e_{i,c}+v_{i,h}+w_{i}e_{i,y}\\ v_{0,h}e_{i,s}-w_{i}e_{i,x}\\ -v_{0,z}+v_{i,z}\\ w_{i}-w_{0}-e_{i,c}w_{i}+e_{i,c}w_{0}\\ e_{i,s}w_{i}-e_{i,s}w_{0}\end{bmatrix}.$ (34)
Note that in Eq.(34), we use a fact that Eq.(35) hold
$\dot{\textbf{p}}^{\mathcal{O}_{0}}_{i}(t)&=\begin{bmatrix}\cos\phi_{\Sigma_{i}}^{\mathcal{O}_{0}}(t),&0\\ \sin\phi_{\Sigma_{i}}^{\mathcal{O}_{0}}(t),&0\\ 0,&1\end{bmatrix}\begin{bmatrix}v_{i,h}\\ v_{i,z}\end{bmatrix}\\ \dot{\phi}_{\Sigma_{i}}^{\mathcal{O}_{0}}(t)&=w_{i}.\end{aligned}\right.$ (35)
With the control law (24), the derivative of $\textbf{e}_{i}$ is,
$\dot{\textbf{e}}_{i}$ $=F(\textbf{e}_{i})+G(\textbf{e}_{i},\tilde{\textbf{e}}_{i}),$ (36)
where $F(\textbf{e}_{i})$ and $G(\textbf{e}_{i},\tilde{\textbf{e}}_{i})$ in (36) are defined as
$F(\textbf{e}_{i})$ $=\begin{bmatrix}-k_{1}e_{i,x}+(1+k_{2})w_{0}e_{i,y}-k_{3}e_{i,s}+v_{0}e_{i,c}\\ v_{0}e_{i,s}-w_{0}e_{i,x}-k_{3}e_{i,s}e_{i,x}\\ -k_{4}e_{i,z}\\ -k_{3}e_{i,s}+k_{3}e_{i,c}e_{i,s}\\ -k_{3}e_{i,s}^{2}\end{bmatrix}$
$G(\textbf{e}_{i},\tilde{\textbf{e}}_{i})$ $=\begin{bmatrix}-k_{1}\tilde{e}_{i,x}-k_{3}\tilde{e}_{i,s}e_{i,y}\\ k_{3}\tilde{e}_{i,s}e_{i,x}\\ -k_{4}\tilde{e}_{i,z}\\ -k_{3}(1-e_{i,c})\tilde{e}_{i,s}\\ -k_{3}e_{i,s}\tilde{e}_{i,s}\end{bmatrix}.$
where $\tilde{e}_{i,x}=\hat{e}_{i,x}-e_{i,x}$, $\tilde{e}_{i,y}=\hat{e}_{i,y}-e_{i,y}$, $\tilde{e}_{i,z}=\hat{e}_{i,z}-e_{i,z}$, $\tilde{e}_{i,c}=\hat{e}_{i,c}-(1-\cos\theta_{\Sigma_{i}}^{\Sigma_{0}})$ and $\tilde{e}_{i,s}=\hat{e}_{i,s}-\sin\theta_{\Sigma_{i}}^{\Sigma_{0}}$ and the estimation error is defined as,
$\tilde{\textbf{e}}_{i}=[\tilde{e}_{i,x},\tilde{e}_{i,y},\tilde{e}_{i,c},\tilde{e}_{i,s}].$ (37)
According to Theorem 2, under Assumption 1 and 2, it is easy to know that the estimated error $\tilde{\textbf{e}}_{i}$ converge to zero asymptotically. According to Proposition 1 in [1], under Assumption 2 and 3, the nominal tracking error system $\dot{\textbf{e}}_{i}=F(\textbf{e}_{i})$ converge to zero exponentially uniformly. Note that the boundness assumption on the leader velocity $v_{0}$ and $w_{0}>0$ are to ensure the uniformly exponential convergence of the nominal tracking error system $\dot{\textbf{e}}_{i}=F(\textbf{e}_{i})$. One can see $G(\textbf{e}_{i},\tilde{\textbf{e}}_{i})$ satisfies
$\lVert G(\textbf{e}_{i},\tilde{\textbf{e}}_{i})\rVert\leq\max\{k_{1},k_{3}\}\lVert\textbf{e}_{i}\rVert\lVert\tilde{\textbf{e}}_{i}\rVert+k_{3}\lVert\textbf{e}_{i}\rVert.$ (38)
According to Lemma 1, the formation tracking error $\textbf{e}_{i}$ converge asymptotically.
-D Complete observability analysis of relative localization
Taking robot $i$ and its neighbor robot $j$ as an example, we discuss several cases where matrix $S_{ij}$ is not full rank and analyze which states cannot be observed.
When robot $i$ remains stationary all the time, i.e. $v_{i}^{\mathcal{O}_{i}}\equiv 0$, both the relative position and the relative yaw is unobservable. If robot $i$ remains stationary all the time, then $\textbf{u}^{\mathcal{O}_{i}}_{i,h}(t_{k})=\textbf{0}$ and $\textbf{p}^{\mathcal{O}_{i}}_{i,h}(t_{k})=\textbf{0}$, then $\mathbf{\Psi}_{ij}(t_{k})=[\textbf{0}_{3\times 1},*_{2\times 1},\textbf{0}_{2\times 1}]^{T}$ and $S_{ij}=\begin{bmatrix}\textbf{0}_{3\times 3},&\textbf{0}_{2\times 2},&\textbf{0}_{2\times 2}\\ \textbf{0}_{3\times 3},&\textbf{*}_{2\times 2},&\textbf{0}_{2\times 2}\\ \textbf{0}_{3\times 3},&\textbf{0}_{2\times 2},&\textbf{0}_{2\times 2}\end{bmatrix}$. The data matrix block corresponding to the relative position and the sine and cosine of and relative yaw angle is always 0, indicating that they are unobservable.
When the neighbor robot $j$ remains stationary all the time, i.e. $v_{j}^{\mathcal{O}_{j}}\equiv 0$, at least the relative yaw is unobservable. One can see that if robot $j$ remains stationary all the time, $\textbf{u}^{\mathcal{O}_{j}}_{j,h}(t_{k})=\textbf{0}$ and $\textbf{p}^{\mathcal{O}_{j}}_{j,h}(t_{k})=\textbf{0}$, then $S_{ij}=\begin{bmatrix}\textbf{*}_{5\times 5},&\textbf{*}_{5\times 2}\\ \textbf{*}_{2\times 5},&\textbf{0}_{2\times 2}\end{bmatrix}$. The data matrix block corresponding to the sine and cosine of and relative yaw angle is always 0, indicating that they are unobservable.
TABLE I: List of main symbols used in this article
Symbols Description
$t_{k}$pbox10cmUWB & IO measurement time instant for robot $i$.
$\mathcal{O}_{i}$pbox10cmThe odometry frame of robot $i$.
$\Sigma_{i}$pbox10cmReal-time IMU body frame of robot $i$.
$\textbf{p}^{\mathcal{O}_{i}}_{i}$pbox10cmDisplacement of robot $i$ in frame $\mathcal{O}_{i}$ from $t_{0}$ to $t_{k}$.
$\phi_{\Sigma_{i}}^{\mathcal{O}_{i}}(t)$pbox10cmAngular displacement of robot $i$ in frame $\mathcal{O}_{i}$ from $t_{0}$ to $t_{k}$.
$\textbf{u}^{\mathcal{O}_{i}}_{i}(t_{k})$pbox10cmDisplacement of robot $i$ in frame $\mathcal{O}_{i}$ from $t_{k}$ to $t_{k+1}$
$d_{ij}(t_{k})$pbox10cmDistance between robots $i$ and $j$.
$y_{ij}$pbox10cmAuxiliary signal which records innovation.
$\hat{\textbf{p}}^{\mathcal{O}_{i}}_{ij,0}$pbox10cm Estimate of initial relative position between robot $i$ and its
neighbor robot $j$ in frame $\mathcal{O}_{i}$.
$\hat{\textbf{R}}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}$pbox10cm Estimate of initial rotation matrix between frame $\mathcal{O}_{i}$ and $\mathcal{O}_{j}$.
$\hat{\textbf{q}}^{\mathcal{O}_{i}}_{i0}$pbox10cm Estimate of initial relative position between robot $i$ and leader
robot 00 in frame $\mathcal{O}_{i}$.
$\hat{\textbf{Q}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}$pbox10cm Estimate of rotation matrix between Frame $\mathcal{O}_{i}$ and $\mathcal{O}_{0}$.
$v_{i}$pbox10cmVelocity control command for robot $i$.
[…]
When the velocity of robot $i$ and its neighbor robot $j$ remain constant, i.e. $v_{i}^{\mathcal{O}_{i}}\equiv v_{i,0}$ and $v_{j}^{\mathcal{O}_{j}}\equiv v_{j,0}$, then the relative position is unobservable. When the velocity of robot $i$ and its neighbor robot $j$ remain constant, one can see that $\textbf{u}^{\mathcal{O}_{i}}_{i,h}(t_{k})^{T}$ and $u^{\mathcal{O}_{i}}_{i,z}(t_{k})-u^{\mathcal{O}_{j}}_{j,z}(t_{k})$ are constant. As a result $S_{ij}=\begin{bmatrix}\textbf{A}_{3\times 3},&\textbf{*}_{3\times 4}\\ \textbf{*}_{4\times 3},&\textbf{*}_{4\times 4}\end{bmatrix}$, where $rank(A)=1$, then the relative position is unobservable.
Appendix A Technical details
A-A Notation list of the proposed method
For ease of reading, in this section, we provide a table of the main symbols that appear in this article in the following Table I.
A-B Pseudocode of the proposed method
The pseudocode of the proposed data-driven localization and formation control method is shown in 1.
Input: UWB and IO measurement. Neighbor IO measurement. Neighbor estimator $\hat{\textbf{q}}^{\mathcal{O}_{i}}_{i0}$ and $\hat{\textbf{Q}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}$.
Output: Control command $v_{i,h}$, $v_{i,z}$ and $w_{i}$
1 while $k+1$-th measurement time instant do
2       Localization enhancement flag = false ;
3       for neighbor $j$ of robot $i$ do
4             Outiler measurment data detetion according to Algorithm 2 ;
5             if measurement data is not outiler then
6                   Collect data according to (10) ;
7                   Update relative pose estimator with (13) ;
8                   if data matrix $S_{ij}$ is not full rank then
9                         Localization enhancement flag = True;
10
11
12      Update cooperative pose estimator with (15) and (16) ;
13       if localization enhancement flag = True then
14             Calculate control command with (23)
15      else
16             Calculate control command with (24
17
Algorithm 1 Localization and control algorithm
In order to reduce the impact of measurement outliers on the localization method, an outlier detection algorithm is proposed in Algorithm 2. In Algorithm 2, we maintain a real-time updated data queue $Q_{ij,\mathrm{judge}}$ (lines 15-17 of the algorithm), with a maximum length of $n_{\mathrm{judge},\mathrm{max}}$. The core idea of the judgment is that the absolute value of the difference between the current measured distance $d_{ij}[k]$ and any distance $d_{ij}[m]$ in the queue should be less than the sum of the robot’s displacement between the two measurements, i.e. $\lVert z_{i}^{\mathcal{O}_{i}}[k]-z_{i}^{\mathcal{O}_{i}}[m]\rVert+\lVert z_{j}^{\mathcal{O}_{j}}[k]-z_{j}^{\mathcal{O}_{j}}[m]\rVert$. This is actually a triangular inequality, as stated in lines 5-9 of the algorithm. If this condition is not met, then there is a problem with the current measurement and it should not be used. To avoid randomness, we set a threshold $\eta_{\mathrm{judge}}$. When the percentage of measurements in the queue that exceeds the threshold does not meet this condition i.e. $n_{\mathrm{outlier}}/n_{\mathrm{judge}}>\eta_{\mathrm{judge}}$, it is considered that the current measurement is incorrect, that is an outlier (lines 10-14 of the algorithm). To reduce the impact of odometer drift, we update the queue with the latest measurement data (lines 15-17 in the algorithm). $n_{\mathrm{judge}}=20$ and $\eta_{\mathrm{judge}}=0.5$.
Input: current data triplet $\mathcal{D}_{ij}[k]=\{d_{ij}[k],z_{i}^{\mathcal{O}_{i}}[k],z_{j}^{\mathcal{O}_{j}}[k]\}$. maximum data queue size $n_{\mathrm{judge},\mathrm{max}}$. Judgment threshold $\eta_{\mathrm{judge}}$.
Output: Outlier detection flag $f_{ij,\mathrm{outlier}}$
1 Judgment data queue $Q_{ij,\mathrm{judge}}$leftarrow empty set $\phi$;
2 while $k$-th measurement time instant do
3       Current data triplet $\mathcal{D}_{ij}[k]$leftarrow UWB & IO measurement;
4       Outlier vote number $n_{\mathrm{outlier}}$leftarrow 0;
5       for data triplet $\mathcal{D}_{ij}[m]\in Q_{\mathrm{judge}}$ do
6             if $|d_{ij}[k]-d_{ij}[m]|\geq\lVert z_{i}^{\mathcal{O}_{i}}[k]-z_{i}^{\mathcal{O}_{i}}[m]\rVert+\lVert z_{j}^{\mathcal{O}_{j}}[k]-z_{j}^{\mathcal{O}_{j}}[m]\rVert$ then
7                   $n_{\mathrm{outlier}}$leftarrow $n_{\mathrm{outlier}}+1$;
8
9
10      Data queue $n_{\mathrm{judge}}$leftarrow data triplets number in $Q_{ij,\mathrm{judge}}$;
11       Outlier detection flag initialize $f_{ij,\mathrm{outlier}}$leftarrow 1;
12       if $n_{\mathrm{outlier}}/n_{\mathrm{judge}}>\eta_{\mathrm{judge}}$ then
13             $f_{ij,\mathrm{outlier}}$leftarrow 0
14      if $f_{ij,\mathrm{outlier}}=1$ then
15             Update $Q_{ij,\mathrm{judge}}$ with the current Measurement data triplet $\mathcal{D}_{ij}[k]$
16
Algorithm 2 UWB outlier detection algorithm
Figure 16: Motion trajectory of the proposed method and method in [14]. A: Proposed method. B: method in [14].
Figure 17: Measurement and communication topology with loop
Appendix B Supplemental simulation and experiment results
B-A Discussion on parameter setting
We explain in detail all the parameters involved in the method proposed in this article and their corresponding meanings below. In addition, for all our simulation and hardware experimental scenarios, we have discussed the parameter selection method and the sensitivity of the proposed method to the parameters.
For the parameters adopted in the cooperative localization and formation control:
The parameters $w_{i}$ and $r_{i}$ denote the angular velocity and the radius of the horizontal circular motion to achieve localization enhancement for each robot. Considering that the accuracy of the UWB module used in the experiment is 0.1m, in order to reflect the significant changes in the relative distance during the motion, it is recommended that the sum of the circular motion diameters of the two neighboring robots be more than 10 times the accuracy error. In our experiment, the radius $r_{i}$ is set to 0.3m for all robots, which can reduce the relative measurement error to a certain extent. In our response to your comment 4, we analyzed some situations where the state cannot be fully observed, and the conclusion can be summarized as follows: in order to achieve effective relative pose estimation, there needs to be a certain relative motion between neighboring robots, that is, there must be differences in the robot’s motion speed and angular velocity. Taking this as a guide, it is recommended that the speed and angular velocity of the neighbor localization enhancement motion differ by 0.5 to 2 times. Considering that the maximum linear velocity of the Turtlebot3 robot we used is 0.22m/s and the maximum angular velocity is 2.5rad/s, with a circular motion radius of 0.3m, we set the angular velocities of the neighboring robots to 0.6rad/s and 0.2rad/s respectively, which can effectively achieve relative motion between neighboring agents without exceeding the maximum linear velocity of the agents.
Similarly, the parameter $c_{i,v}$ represents the amplitude and phase of longitudinal motion. To achieve the relative motion of neighboring drones, the selection of this parameter is related to the actual flight altitude. If the flight altitude is low, $c_{i,v}$ cannot be too large to prevent the drone from touching the ground during descent. Specifically, the maximum downward motion amplitude of the drone can be calculated. It is recommended to widen the gradient of the $c_{i,v}$ parameter of neighboring drones while ensuring that the aircraft does not touch the ground. For example, in our drone experiment, we set it to 0.1 and 0.3, respectively.
Figure 18: Pose estimation error and formation control error for each robot in the swarm. A: Estimation error. B: Formation control error.
Figure 19: Velocity and angle velocity for each robot in the swarm.
The control gain $k_{1}$, $k_{2}$, $k_{3}$ and $k_{4}$ only needs to be greater than 0 to ensure asymptotic convergence of the control error. Generally speaking, when the gain value is larger, the algorithm converges faster. However, in computer control systems, due to the discrete control instructions (such as UAV control instructions generally ranging from 20 to 50HZ), divergence or numerical instability may occur when the gain is too large. After multiple experimental verifications, we recommend that the gain $k_{1}$, $k_{2}$, $k_{3}$ and $k_{4}$ be between 1 and 2.
B-B Motion trajectory of comparative simulation
When there are 5 robots in the swarm, based on our proposed method and the existing PE condition based method [14], the motion trajectory of the robots is shown in Fig. 16, which clearly demonstrates the advantages of the proposed data-driven method in improving control performance.
B-C Formation control for large scale swarm
50 robots formed a random topology based on their initial positions, which included loops. The topology diagram is shown in Figure. 17. The parameters adopted are the same with those in the multi-robot case in Section IV-A. The relative pose estimation error and formation control tracking error of all individuals in the swarm for themselves and the leader are shown in Fig. 18(A) and Fig. 18(B), and the velocity and angular velocity of all individuals in the swarm are shown in Fig. 19. It can be seen that the relative pose estimation error and tracking control error of the final individual converge, and the velocity and angular velocity tend to be a consistent value. (i.e., the velocity and angular velocity of the leader).
B-D Noise robustness for large scale swarm
Figure 20: Pose estimation and tracking control error under different noise level. A: Estimation error. B: Tracking control error.
We consider the cooperative estimation error and formation control error of a swarm consisting of 50 robots under different noise conditions. The parameters adopted are the same as those in the multi-robot case in Section IV-A. Specifically, under each noise condition, the algorithm is executed under 100 different initial conditions, and the statistical results of all robot estimation and control errors obtained in each simulation experiment are recorded as shown in Figure 20. It can be seen that under normal noise levels (0.1m), the algorithm has good consistency with noise, but its performance decreases as the noise level increases. In addition, due to the large size of the swarm, the error of cooperative localization may propagate with the increase of topology layers, but this can be solved through topology layering, adding leaders, and other methods, which can be discussed in future work.
B-E Sensor highly inaccurate case
Figure 21: Accuracy of localization algorithm with and without outlier detection algorithm for sensors at different sensor failure rates.
On the basis of simulation in Section IV-B, we also simulated sensor damage during 100 simulations, which means that after a random moment in the simulation until the end of the simulation time, the sensor outputs measurement values with huge noise (5 as variance) with a 0.95 probability. The parameters adopted are the same as those in the two-robot case in Section IV-A. Under such unfavorable conditions, the accuracy of the localization method is shown in Figure 21. It can be seen that the outlier detection algorithm can improve the accuracy of localization performance to a certain extent.
B-F Formation control with different initial state
The snapshot of experiments starting from three different initial conditions in ten sets of experiments is shown in Figure. 22.
Figure 22: Snapshot of experiments starting from three different initial conditions in ten sets of experiments.